A department store sells a 4-pack of AAA batteries for $3.00 and an 8-pack of AAA batteries for $5.00. The cost of one battery from the 4-pack is what percent more than the cost of one battery from the 8-pack, before sales tax is applied?

Answer by stanbon(75887)

A department store sells a 4-pack of AAA batteries for $3.00 and an 8-pack of AAA batteries for $5.00. The cost of one battery from the 4-pack is what percent more than the cost of one battery from the 8-pack, before sales tax is applied?
-------------
Each battery from the 4-pack casts 300/4 = 75 cents
----
Each battery from the 8-pack costs 500/8 = 62.5 cents
----
Quantity difference is 75-62.5 = 12.5 cents
---
Percent difference = 12.5/62.5 = 0.20 = 20%
================================================
Cheers,
Stan H.
==========